OpenStudy (anonymous):

graph linear equation by solving for y first -2x+3y=-3

OpenStudy (anonymous):

Move the variables around for slope intercept form. y = mx + b 3y = 2x - 3 Then just divide the 3 to the other side y = (2/3)x - 1 Y-intercept is -1, slope is 2/3

OpenStudy (anonymous):

how would you put that in graph

OpenStudy (anonymous):

As you can see, because the y-intercept is -1, it will cross the -1 on the y-axis. As for the slope, which is "m" in the equation "y=mx+b", it's 2/3. Before you graph the slope, first make sure you have the y-intercept graphed. Now, to graph the slope, just remember the phrase "rise/run". If 2/3 is the slope, then you start from the y-intercept, and you rise 2, then run 3. In other words, from (0, -1), you go up 2, then to the right 3, and plot a point there. Now keep repeating the rise/run process from the points you plot and once you've plotted down 2 or 3 points, you should be able to draw a straight line between the points. Now, you have a graph! :)

OpenStudy (anonymous):

thank you